# Sampling the space of solutions of an artificial neural network

### v1

@@ B.1 Replica approach
The computation proceeds as usual introducing replicas via the identity $(Z_{\mathcal{D}}^{r})^{-1}=\lim\limits_{n\to 0}(Z_{\mathcal{D}}^{r})^{n-1}$ with $r=1,2$. We have (denoting with $r,s=1,2$ the index that runs over the real replicas $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$)
$\begin{split}E(\gamma)&=\int\prod_{arl\mu}\frac{d\lambda_{lr}^{\mu a}d\hat{\lambda}_{lr}^{\mu a}}{2\pi}e^{i\lambda_{lr}^{\mu a}\hat{\lambda}_{lr}^{\mu a}}\int\prod_{ar}d\boldsymbol{w}^{ra}\,\prod_{ra\mu}e^{-\beta\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(\lambda^{\mu a}_{lr}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\frac{\gamma\lambda_{l1}^{\mu 1}+(1-\gamma)\lambda_{l2}^{\mu 1}}{c_{\gamma}}\right)\right]\\ &\times e^{-\frac{\lambda_{1}}{2}\sum_{lia}(w_{li}^{ar=1})^{2}-\frac{\lambda_{2}}{2}\sum_{lia}(w_{li}^{ar=2})^{2}}\prod_{li\mu}\mathbb{E}_{\xi_{li}^{\mu}}e^{-i\frac{\xi_{li}^{\mu}}{\sqrt{N}}\sum_{ar}w_{li}^{ar}\hat{\lambda}_{lr}^{\mu a}}\\ &=\int\prod_{ar,bs,l}\frac{dq_{rsl}^{ab}d\hat{q}_{rsl}^{ab}}{2\pi}e^{-N\sum_{ar,bsl}q_{rsl}^{ab}\hat{q}_{rsl}^{ab}-\frac{\lambda_{1}}{2}\sum_{la}q^{aa}_{11l}-\frac{\lambda_{2}}{2}\sum_{rla}q^{aa}_{22l}+NG_{S}+(N\alpha-1)G_{E}}\\ &\times\int\prod_{arl}\frac{d\lambda_{lr}^{a}d\hat{\lambda}_{lr}^{a}}{2\pi}\,e^{i\lambda_{lr}^{a}\hat{\lambda}_{lr}^{a}}\prod_{ra}e^{-\beta\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(\lambda^{\mu a}_{lr}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\frac{\gamma\lambda_{l1}^{\mu 1}+(1-\gamma)\lambda_{l2}^{\mu 1}}{c_{\gamma}}\right)\right]e^{-\frac{1}{2}\sum_{ar,bs}q^{ab}_{rs}\hat{\lambda}_{r}^{a}\hat{\lambda}_{s}^{b}}\end{split}$ (40)
where we remind that $c_{\gamma}$ is the the quantity that appears in equation (24). We have also introduced the usual entropic and energetic terms Baldassi et al. (2019) as
$\begin{split}G_{S}&\equiv\ln\int\prod_{ar}dw^{ra}_{l}\,e^{\sum_{ar,bsl}\hat{q}^{ab}_{rsl}w_{l}^{ra}w_{l}^{sb}}\\ G_{E}&\equiv\ln\int\prod_{arl}\frac{d\lambda_{rl}^{a}d\hat{\lambda}_{rl}^{a}}{2\pi}\,e^{-\beta\sum_{ra}\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\lambda^{a}_{rl}\right)\right)}\,e^{i\sum_{rla}\lambda_{rl}^{a}\hat{\lambda}_{rl}^{a}-\frac{1}{2}\sum_{ar,bsl}q^{ab}_{rsl}\hat{\lambda}_{rl}^{a}\hat{\lambda}_{sl}^{b}}\,.\end{split}$ (41)
@@ B.2 Replica Symmetric ansatz
Notice that we called by $q_{1}$ and $q_{2}$ the typical overlap between solutions extracted from the distribution of the endpoint $\gamma=1$ and $\gamma=0$ respectively. The overlap $p$ represents the typical overlap between the two endpoints. A similar ansatz is imposed over the conjugated order parameters $\hat{q}_{rsl}^{ab}$. However in the $n\to 0$ limit the conjugated order parameters will not appear explicitly in the expression of $E(\gamma)$. We need to express decompose the term
$\begin{split}-\frac{1}{2}\sum_{abrsl}q_{rsl}^{ab}\hat{\lambda}_{lr}^{a}\hat{\lambda}_{ls}^{b}&=-\frac{1}{2}\sum_{r}(Q-q_{r})\sum_{al}\left(\hat{\lambda}_{rl}^{a}\right)^{2}-\frac{1}{2}\sum_{rs}t_{rs}\sum_{abl}\hat{\lambda}_{rl}^{a}\hat{\lambda}_{sl}^{b}\\ &=-\frac{1}{2}\sum_{r}(Q-q_{r})\sum_{al}\left(\hat{\lambda}_{rl}^{a}\right)^{2}-\frac{1}{2}\sum_{rl}\left(\sum_{as}\mathcal{T}_{rs}\hat{\lambda}^{a}_{sl}\right)^{2}\end{split}$ (43)
@@
where $\mathcal{T}_{rs}$ is $(r,s)$ the element of the square root of the matrix $t_{rs}$ defined in (42b). The computation proceeds in a standard way by using a Hubbard-Stratonovich transformation and integrating over $\hat{\lambda}_{rl}^{a}$. We get
@@
$\begin{split}E(\gamma)&=\int\prod_{arl}\frac{d\lambda_{rl}^{a}d\hat{\lambda}_{rl}^{a}}{2\pi}\,e^{-\beta\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(\lambda^{a}_{lr}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\frac{\gamma\lambda_{l1}^{1}+(1-\gamma)\lambda_{l2}^{1}}{c_{\gamma}}\right)\right]\\ &\times e^{i\sum_{arl}\lambda_{rl}^{a}\hat{\lambda}_{rl}^{a}-\frac{1}{2}\sum_{r}(Q-q_{r})\sum_{al}\left(\hat{\lambda}_{rl}^{a}\right)^{2}-\frac{1}{2}\sum_{rl}\left(\sum_{as}\mathcal{T}_{rs}\hat{\lambda}^{a}_{sl}\right)^{2}}\\ &=\int\!\prod_{rl}Dx_{rl}\frac{\int\prod_{rl}\frac{d\lambda_{rl}d\hat{\lambda}_{rl}}{2\pi}\,e^{i\hat{\lambda}_{rl}\left(\lambda_{rl}+\sum_{s}\mathcal{T}_{rs}x_{sl}\right)-\frac{1}{2}(Q-q_{r})\hat{\lambda}_{rl}^{2}-\beta\sum_{r}\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(\lambda_{lr}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\frac{\gamma\lambda_{l1}+(1-\gamma)\lambda_{l2}}{c_{\gamma}}\right)\right]}{\prod_{r}\int\prod_{l}\frac{d\lambda_{l}d\hat{\lambda}_{l}}{2\pi}\,e^{-\beta\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(\lambda_{l}\right)\right)}\,e^{i\sum_{l}\hat{\lambda}_{l}\left(\lambda_{l}+\sum_{s}\mathcal{T}_{sr}\,x_{rl}\right)-\frac{1}{2}\sum_{r}(Q-q_{r})\sum_{l}\hat{\lambda}_{l}^{2}}}\\ &=\int\!\prod_{rl}Dx_{rl}\frac{\int\prod_{rl}D\lambda_{rl}\,e^{-\beta\sum_{r}\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(\sqrt{Q-q_{r}}\lambda_{lr}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}\lambda_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)\right]}{\prod_{r}\int\prod_{l}D\lambda_{l}\,e^{-\beta\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(\sqrt{Q-q_{r}}\lambda_{l}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)\right)}}\end{split}$ (44)
@@ B.3 Large $K$ limit
$\begin{split}I&\equiv\int\prod_{rl}D\lambda_{rl}\,e^{-\beta\sum_{r}\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(\sqrt{Q-q_{r}}\lambda_{lr}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}\lambda_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)\right]\\ &=\int\prod_{r}\frac{dh_{r}d\hat{h}_{r}}{2\pi}\,e^{i\sum_{r}h_{r}\hat{h}_{r}-\beta\sum_{r}\ell_{r}\left(h_{r}\right)}\int\frac{dvd\hat{v}}{2\pi}e^{i\hat{v}v}\,\tilde{\ell}\left[v\right]\\ &\times\int\prod_{rl}D\lambda_{rl}\,e^{-\frac{i}{\sqrt{K}}\sum_{r}\hat{h}_{r}\sum_{l}c_{l}\varphi\left(\sqrt{Q-q_{r}}\lambda_{lr}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)-\frac{i\hat{v}}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}\lambda_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)}\end{split}$ (45)
Expanding the exponential up to second order, averaging over $\lambda_{lr}$ and re-exponentiating we get
$\begin{split}&\int\prod_{rl}D\lambda_{rl}\,e^{-\frac{i}{\sqrt{K}}\sum_{r}\hat{h}_{r}\sum_{l}c_{l}\varphi\left(\sqrt{Q-q_{r}}\lambda_{lr}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)-\frac{i\hat{v}}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}\lambda_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)}\\ &\simeq e^{-i\sum_{r}\hat{h}_{r}M_{r}^{(0)}-i\hat{v}N^{(0)}-\frac{1}{2}\sum_{r}\Delta_{r}^{(0)}\hat{h}_{r}^{2}-\frac{\Xi^{(0)}}{2}\hat{v}^{2}-\hat{v}\sum_{r}\hat{h}_{r}\Omega_{r}^{(0)}}\end{split}$ (46)
@@
$\varphi_{rl}(\lambda_{r},x)$ $\equiv\varphi\left(\sqrt{Q-q_{r}}\lambda_{lr}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)$ (47a)
$\tilde{\varphi}_{l}(\gamma;\{\lambda_{r}\}_{r=1}^{y},x)$ $\equiv\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}\lambda_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)$ (47b)
@@
$M^{(0)}_{r}$ $\equiv\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\left\langle\varphi\left(\sqrt{Q-q_{r}}\lambda_{r}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)\right\rangle_{\lambda_{r}}\equiv\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\left\langle\varphi_{rl}\right\rangle_{\lambda}$ (48a)
$N^{(0)}$ $\equiv\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\left\langle\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}\lambda_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)\right\rangle_{\{\lambda_{r}\}_{r=1}^{y}}\equiv\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\left\langle\tilde{\varphi}_{l}\right\rangle_{\lambda}$ (48b)
$\Delta^{(0)}_{r}$ $\equiv\frac{1}{K}\sum_{l}c_{l}^{2}\,\left[\langle\varphi^{2}_{rl}\rangle_{\lambda}-\langle\varphi_{rl}\rangle^{2}_{\lambda}\right]$ (48c)
$\Xi^{(0)}$ $\equiv\frac{1}{K}\sum_{l}c_{l}^{2}\left[\langle\tilde{\varphi}^{2}_{l}\rangle_{\lambda}-\langle\tilde{\varphi}_{l}\rangle^{2}_{\lambda}\right]$ (48d)
$\Omega_{r}^{(0)}$ $\equiv\frac{1}{K}\sum_{l}c_{l}^{2}\left[\langle\tilde{\varphi}_{l}\varphi_{lr}\rangle_{\lambda}-\langle\tilde{\varphi}_{l}\rangle_{\lambda}\langle\varphi_{lr}\rangle_{\lambda_{r}}\right]$ (48e)
Notice that $\Delta^{(0)}_{rs}=\frac{1}{K}\sum_{l}c_{l}^{2}\,\left[\langle\varphi_{rl}\varphi_{sl}\rangle_{\lambda}-\langle\varphi_{rl}\rangle_{\lambda}\langle\varphi_{sl}\rangle_{\lambda}\right]=0$ if $r\neq s$ since the Gaussian variable $\lambda_{lr}$ is independent of $\lambda_{ls}$. We then inserting (46) back into (45) and perform the integrals over $\hat{h}_{r}$ and $\hat{v}$; in general the integral is of the form
$\begin{split}&\int\prod_{r}\frac{dh_{r}d\hat{h}_{r}}{2\pi}\frac{dvd\hat{v}}{2\pi}e^{i\hat{h}_{r}(h_{r}-M_{r})+i\hat{v}(v-N)-\frac{1}{2}\sum_{rs}\Delta_{rs}\hat{h}_{r}\hat{h}_{s}-\frac{\Xi}{2}\hat{v}^{2}-\hat{v}\sum_{r}\Omega_{r}\hat{h}_{r}}f(\{h_{r}\},v)\\ &\simeq\int\prod_{r}\frac{dh_{r}d\hat{h}_{r}}{2\pi}Dv\,e^{i\sum_{r}\hat{h}_{r}(h_{r}-M_{r}-\frac{v}{\sqrt{\Xi}}\Omega_{r})-\frac{1}{2}\sum_{rs}\left(\Delta_{rs}-\frac{\Omega_{r}\Omega_{s}}{\Xi}\right)\hat{h}_{r}\hat{h}_{s}}f(\{h_{r}\},N+\sqrt{\Xi}v)\\ &=\int Dv\prod_{r}Dh_{r}\,f\left(\left\{M_{r}+\frac{\Omega_{r}}{\sqrt{\Xi}}\,v+\sum_{s}\left(\sqrt{A}\right)_{rs}h_{s}\right\},N+\sqrt{\Xi}v\right)\end{split}$ (49)
with $A_{rs}\equiv\Delta_{rs}-\frac{\Omega_{r}\Omega_{s}}{\Xi}$. Specializing this identity to our case, i.e. using $f(\left\{h_{r}\right\},v)=e^{-\beta\sum_{r}\ell_{r}(h_{r})}\,\widetilde{\ell}(v)$ we get

### v2

@@ B.1 Replica approach
The computation proceeds as usual introducing replicas via the identity $(Z_{\mathcal{D}}^{r})^{-1}=\lim\limits_{n\to 0}(Z_{\mathcal{D}}^{r})^{n-1}$ with $r=1,2$. In the following we avoid explicating the range of each sum and product for simplicity and to lighten up notation. We however specify here that the indexes $r,s=1,2$ will run over the real replicas $\boldsymbol{w}^{1}$ and $\boldsymbol{w}^{2}$, the index $\mu=1,\dots,P$ will denote the data index, the indexes $a,b=1,\dots,n$ will run over the $n$ replicas of the system, the index $l=1,\dots,K$ will run over the $K$ hidden units and finally $i=1,\dots,N/K$ will be the index running over the weights in single branch of the tree network.
We have
$\begin{split}E(\gamma)&=\int\prod_{arl\mu}\frac{dv_{lr}^{\mu a}d\hat{v}_{lr}^{\mu a}}{2\pi}e^{iv_{lr}^{\mu a}\hat{v}_{lr}^{\mu a}}\int\prod_{ar}d\boldsymbol{w}^{ra}\,\prod_{ra\mu}e^{-\beta\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(v^{\mu a}_{lr}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\frac{\gamma v_{l1}^{\mu_{\star}1}+(1-\gamma)v_{l2}^{\mu_{\star}1}}{c_{\gamma}}\right)\right]\\ &\times e^{-\frac{\lambda_{1}}{2}\sum_{lia}(w_{li}^{ar=1})^{2}-\frac{\lambda_{2}}{2}\sum_{lia}(w_{li}^{ar=2})^{2}}\prod_{li\mu}\mathbb{E}_{\xi_{li}^{\mu}}e^{-i\frac{\xi_{li}^{\mu}}{\sqrt{N}}\sum_{ar}w_{li}^{ar}\hat{v}_{lr}^{\mu a}}\\ &=\int\prod_{ar,bs,l}\frac{dq_{rsl}^{ab}d\hat{q}_{rsl}^{ab}}{2\pi}e^{-N\sum_{ar,bsl}q_{rsl}^{ab}\hat{q}_{rsl}^{ab}-\frac{\lambda_{1}}{2}\sum_{la}q^{aa}_{11l}-\frac{\lambda_{2}}{2}\sum_{rla}q^{aa}_{22l}+NG_{S}+(N\alpha-1)G_{E}}\\ &\times\int\prod_{arl}\frac{dv_{lr}^{a}d\hat{v}_{lr}^{a}}{2\pi}\,e^{iv_{lr}^{a}\hat{v}_{lr}^{a}}\prod_{ra}e^{-\beta\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(v^{a}_{lr}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\frac{\gamma v_{l1}^{1}+(1-\gamma)v_{l2}^{1}}{c_{\gamma}}\right)\right]e^{-\frac{1}{2}\sum_{ar,bsl}q^{ab}_{rsl}\hat{v}_{lr}^{a}\hat{v}_{ls}^{b}}\end{split}$ (41)
where we remind that $c_{\gamma}$ is the the quantity that appears in equation (24). In the second equality of the previous equation we have performed the average over the Gaussian input data and introduced the following overlaps quantities
$q^{ab}_{rsl}\equiv\frac{K}{N}\sum_{i=1}^{N/K}w^{ar}_{li}w^{bs}_{li}$ (42)
via delta functions and their integral representations. This enables the decoupling over the site index $i=1,\dots,N/K$ and the ($P-1$) pattern index $\mu\neq\mu_{\star}$. The last line of (41) refers to the contribution coming from the pattern index $\mu_{\star}$ whose dependence we have dropped for simplicity. We have also introduced the usual entropic and energetic terms Baldassi et al. (2019) as
$\begin{split}G_{S}&\equiv\frac{1}{K}\ln\int\prod_{arl}dw^{ra}_{l}\,e^{\sum_{ar,bsl}\hat{q}^{ab}_{rsl}w_{l}^{ra}w_{l}^{sb}}\\ G_{E}&\equiv\ln\int\prod_{arl}\frac{dv_{rl}^{a}d\hat{v}_{rl}^{a}}{2\pi}\,e^{-\beta\sum_{ra}\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(v^{a}_{rl}\right)\right)}\,e^{i\sum_{rla}v_{rl}^{a}\hat{v}_{rl}^{a}-\frac{1}{2}\sum_{ar,bsl}q^{ab}_{rsl}\hat{v}_{rl}^{a}\hat{v}_{sl}^{b}}\,.\end{split}$ (43)
@@ B.2 Replica Symmetric ansatz
Notice that we called by $q_{1}$ and $q_{2}$ the typical overlap between solutions extracted from the distribution of the endpoint $\gamma=1$ and $\gamma=0$ respectively. The overlap $p$ represents the typical overlap between the two endpoints. A similar ansatz is imposed over the conjugated order parameters $\hat{q}_{rsl}^{ab}$. However in the $n\to 0$ limit the conjugated order parameters will not appear explicitly in the expression of $E(\gamma)$, as the first line after the equal sign in (41) goes to 1 in the small $n$ limit. We need to express decompose the term
$\begin{split}-\frac{1}{2}\sum_{abrsl}q_{rsl}^{ab}\hat{v}_{lr}^{a}\hat{v}_{ls}^{b}&=-\frac{1}{2}\sum_{r}(Q-q_{r})\sum_{al}\left(\hat{v}_{rl}^{a}\right)^{2}-\frac{1}{2}\sum_{rs}t_{rs}\sum_{abl}\hat{v}_{rl}^{a}\hat{v}_{sl}^{b}\\ &=-\frac{1}{2}\sum_{r}(Q-q_{r})\sum_{al}\left(\hat{v}_{rl}^{a}\right)^{2}-\frac{1}{2}\sum_{rl}\left(\sum_{as}\mathcal{T}_{rs}\hat{v}^{a}_{sl}\right)^{2}\end{split}$ (45)
@@
where $\mathcal{T}_{rs}$ is the $(r,s)$ element of the square root of the matrix $t_{rs}$ defined in (44b). The computation proceeds in a standard way by using a Hubbard-Stratonovich transformation and integrating over $\hat{v}_{rl}^{a}$. We get
@@
$\begin{split}E(\gamma)&=\int\prod_{arl}\frac{dv_{rl}^{a}d\hat{v}_{rl}^{a}}{2\pi}\,e^{-\beta\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(v^{a}_{lr}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\frac{\gamma v_{l1}^{1}+(1-\gamma)v_{l2}^{1}}{c_{\gamma}}\right)\right]\\ &\times e^{i\sum_{arl}v_{rl}^{a}\hat{v}_{rl}^{a}-\frac{1}{2}\sum_{r}(Q-q_{r})\sum_{al}\left(\hat{v}_{rl}^{a}\right)^{2}-\frac{1}{2}\sum_{rl}\left(\sum_{as}\mathcal{T}_{rs}\hat{v}^{a}_{sl}\right)^{2}}\\ &=\int\!\prod_{rl}Dx_{rl}\frac{\int\prod_{rl}\frac{dv_{rl}d\hat{v}_{rl}}{2\pi}\,e^{i\hat{v}_{rl}\left(v_{rl}+\sum_{s}\mathcal{T}_{rs}x_{sl}\right)-\frac{1}{2}(Q-q_{r})\hat{v}_{rl}^{2}-\beta\sum_{r}\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(v_{lr}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\frac{\gamma v_{l1}+(1-\gamma)v_{l2}}{c_{\gamma}}\right)\right]}{\prod_{r}\int\prod_{l}\frac{dv_{l}d\hat{v}_{l}}{2\pi}\,e^{-\beta\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(v_{l}\right)\right)}\,e^{i\sum_{l}\hat{v}_{l}\left(v_{l}+\sum_{s}\mathcal{T}_{sr}\,x_{rl}\right)-\frac{1}{2}\sum_{r}(Q-q_{r})\sum_{l}\hat{v}_{l}^{2}}}\\ &=\int\!\prod_{rl}Dx_{rl}\frac{\int\prod_{rl}Dv_{rl}\,e^{-\beta\sum_{r}\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(\sqrt{Q-q_{r}}v_{lr}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}v_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)\right]}{\prod_{r}\int\prod_{l}Dv_{l}\,e^{-\beta\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(\sqrt{Q-q_{r}}v_{l}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)\right)}}\end{split}$ (46)
@@ B.3 Large $K$ limit
$\begin{split}I&\equiv\int\prod_{rl}Dv_{rl}\,e^{-\beta\sum_{r}\ell_{r}\left(\frac{1}{\sqrt{K}}\sum_{l}c_{l}\varphi\left(\sqrt{Q-q_{r}}v_{lr}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)\right)}\,\tilde{\ell}\left[\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}v_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)\right]\\ &=\int\prod_{r}\frac{dh_{r}d\hat{h}_{r}}{2\pi}\,e^{i\sum_{r}h_{r}\hat{h}_{r}-\beta\sum_{r}\ell_{r}\left(h_{r}\right)}\int\frac{d\eta d\hat{\eta}}{2\pi}e^{i\hat{\eta}\eta}\,\tilde{\ell}\left[\eta\right]\\ &\times\int\prod_{rl}Dv_{rl}\,e^{-\frac{i}{\sqrt{K}}\sum_{r}\hat{h}_{r}\sum_{l}c_{l}\varphi\left(\sqrt{Q-q_{r}}v_{lr}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)-\frac{i\hat{\eta}}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}v_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)}\end{split}$ (47)
Expanding the exponential up to second order, averaging over $v_{lr}$ and re-exponentiating we get
$\begin{split}&\int\prod_{rl}Dv_{rl}\,e^{-\frac{i}{\sqrt{K}}\sum_{r}\hat{h}_{r}\sum_{l}c_{l}\varphi\left(\sqrt{Q-q_{r}}v_{lr}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)-\frac{i\hat{\eta}}{\sqrt{K}}\sum_{l}c_{l}\,\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}v_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)}\\ &\simeq e^{-i\sum_{r}\hat{h}_{r}M_{r}^{(0)}-i\hat{\eta}N^{(0)}-\frac{1}{2}\sum_{r}\Delta_{r}^{(0)}\hat{h}_{r}^{2}-\frac{\Xi^{(0)}}{2}\hat{\eta}^{2}-\hat{\eta}\sum_{r}\hat{h}_{r}\Omega_{r}^{(0)}}\end{split}$ (48)
@@
$\varphi_{rl}(v_{r},x)$ $\equiv\varphi\left(\sqrt{Q-q_{r}}v_{lr}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)$ (49a)
$\tilde{\varphi}_{l}(\gamma;\{v_{r}\}_{r=1}^{y},x)$ $\equiv\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}v_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)$ (49b)
@@
$M^{(0)}_{r}$ $\equiv\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\left\langle\varphi\left(\sqrt{Q-q_{r}}v_{r}-\sum_{s}\mathcal{T}_{rs}x_{sl}\right)\right\rangle_{v_{r}}\equiv\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\left\langle\varphi_{rl}\right\rangle_{v}$ (50a)
$N^{(0)}$ $\equiv\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\left\langle\varphi\left(\sum_{r}\gamma_{r}\left(\frac{\sqrt{Q-q_{r}}v_{rl}-\sum_{s}\mathcal{T}_{rs}x_{sl}}{c_{\gamma}}\right)\right)\right\rangle_{\{v_{r}\}_{r=1}^{y}}\equiv\frac{1}{\sqrt{K}}\sum_{l}c_{l}\,\left\langle\tilde{\varphi}_{l}\right\rangle_{v}$ (50b)
$\Delta^{(0)}_{r}$ $\equiv\frac{1}{K}\sum_{l}c_{l}^{2}\,\left[\langle\varphi^{2}_{rl}\rangle_{v}-\langle\varphi_{rl}\rangle^{2}_{v}\right]$ (50c)
$\Xi^{(0)}$ $\equiv\frac{1}{K}\sum_{l}c_{l}^{2}\left[\langle\tilde{\varphi}^{2}_{l}\rangle_{v}-\langle\tilde{\varphi}_{l}\rangle^{2}_{v}\right]$ (50d)
$\Omega_{r}^{(0)}$ $\equiv\frac{1}{K}\sum_{l}c_{l}^{2}\left[\langle\tilde{\varphi}_{l}\varphi_{lr}\rangle_{v}-\langle\tilde{\varphi}_{l}\rangle_{v}\langle\varphi_{lr}\rangle_{v_{r}}\right]$ (50e)
Notice that $\Delta^{(0)}_{rs}=\frac{1}{K}\sum_{l}c_{l}^{2}\,\left[\langle\varphi_{rl}\varphi_{sl}\rangle_{v}-\langle\varphi_{rl}\rangle_{v}\langle\varphi_{sl}\rangle_{v}\right]=0$ if $r\neq s$ since the Gaussian variable $v_{lr}$ is independent of $v_{ls}$. We then inserting (48) back into (47) and perform the integrals over $\hat{h}_{r}$ and $\hat{v}$; in general the integral is of the form
$\begin{split}&\int\prod_{r}\frac{dh_{r}d\hat{h}_{r}}{2\pi}\frac{d\eta d\hat{\eta}}{2\pi}e^{i\hat{h}_{r}(h_{r}-M_{r})+i\hat{\eta}(\eta-N)-\frac{1}{2}\sum_{rs}\Delta_{rs}\hat{h}_{r}\hat{h}_{s}-\frac{\Xi}{2}\hat{\eta}^{2}-\hat{\eta}\sum_{r}\Omega_{r}\hat{h}_{r}}f(\{h_{r}\},\eta)\\ &\simeq\int\prod_{r}\frac{dh_{r}d\hat{h}_{r}}{2\pi}D\eta\,e^{i\sum_{r}\hat{h}_{r}(h_{r}-M_{r}-\frac{\eta}{\sqrt{\Xi}}\Omega_{r})-\frac{1}{2}\sum_{rs}\left(\Delta_{rs}-\frac{\Omega_{r}\Omega_{s}}{\Xi}\right)\hat{h}_{r}\hat{h}_{s}}f(\{h_{r}\},N+\sqrt{\Xi}\eta)\\ &=\int D\eta\prod_{r}Dh_{r}\,f\left(\left\{M_{r}+\frac{\Omega_{r}}{\sqrt{\Xi}}\,\eta+\sum_{s}\left(\sqrt{A}\right)_{rs}h_{s}\right\},N+\sqrt{\Xi}\eta\right)\end{split}$ (51)
with $A_{rs}\equiv\Delta_{rs}-\frac{\Omega_{r}\Omega_{s}}{\Xi}$. Specializing this identity to our case, i.e. using $f(\left\{h_{r}\right\},\eta)=e^{-\beta\sum_{r}\ell_{r}(h_{r})}\,\widetilde{\ell}(\eta)$ we get